# The $\beta$-decay spectrum of Tritiated graphene: combining nuclear quantum mechanics with Density Functional Theory

### v1

@@ Acknowledgments
We are grateful to the whole PTOLEMY collaboration for precious discussions and comments on the manuscript. AE is grateful to José Lorenzana for enlightening discussions on the role of the sudden approximation, and for useful suggestions. GM and VT also thank Paolo Giannozzi for very useful discussions and support in DFT calculations. The work of AC and AE has been supported by Sapienza University through the SEED PNR 2022 funds. GM and VT acknowledge the allocation of HPC resources on LEONARDO-CINECA trough ISCRA B SETE (n. HP10BSTFDD) and through the ICSC (Centro Nazionale di Ricerca HPC, Big Data and Quantum Coputing, grant n. 1491920). This research was also supported by EU under FETPROACT LESGO (Agreement No. 952068).

### v2

@@ Acknowledgments
We are grateful to the whole PTOLEMY collaboration for precious discussions and comments on the manuscript. AE is grateful to JosÊ Lorenzana for enlightening discussions on the role of the sudden approximation, and for useful suggestions. GM and VT also thank Paolo Giannozzi for very useful discussions and support in DFT calculations. The work of AC and AE has been supported by Sapienza University through the SEED PNR 2022 funds. GM and VT acknowledge the allocation of HPC resources on LEONARDO-CINECA trough ISCRA B SETE (n. HP10BSTFDD) and through the ICSC (Centro Nazionale di Ricerca HPC, Big Data and Quantum Computing, grant n. 1491920). This research was also supported by EU under FETPROACT LESGO (Agreement No. 952068).